OpenStudy (anonymous):

Solve the following equation: x2 + 14x + 49 = 0 A. -7 B. {0, 7} C. 7 D. {7, -7}

OpenStudy (anonymous):

Factor it what are the factors of 49?

OpenStudy (anonymous):

that add up to 14?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Yes, so x^2 +7x+7x+49=0 ---> (x+7)(x+7)=0

OpenStudy (anonymous):

Now, solve this and see what your answer is

OpenStudy (anonymous):

So B. ?

OpenStudy (anonymous):

No x+7=0 or x+7=0

OpenStudy (anonymous):

which becomess...

OpenStudy (anonymous):

7

OpenStudy (anonymous):

No... x+7=0 is the same as x=-7

OpenStudy (anonymous):

So A

OpenStudy (anonymous):

Yes